OpenStudy (anonymous):

If F(x) = x² + 5x and G(x) = 2x + 1, find F(5) + G(6).

OpenStudy (anonymous):

63

OpenStudy (anonymous):

F(x) =( 5² + 5(5))+( 2(6) + 1) = (25+25)+(12+1) =50+13 = 63

OpenStudy (anonymous):

\[f(x)=x² + 5x\] \[f(5)=5² + 5.5\] \[f(5) = 50\] \[g(x)=2x +1\] \[g(6)=2.6 +1\] \[g(6)=13\] \[f(5) + g(6) = 63\]